OpenStudy (anonymous):

Solve for x: 3x = 18 im pretty sure its 6

jigglypuff314 (jigglypuff314):

6 is correct! ^_^

OpenStudy (_michaeld22):

You're right, it's 6.

OpenStudy (_michaeld22):

Ahh @jigglypuff314 Beat me to it :P

OpenStudy (anonymous):

2x - 8 = -24 I think its 8 but don't know if its -8

jigglypuff314 (jigglypuff314):

add 8 to both sides 2x - 8 = -24 +8 +8 so 2x = -16 divide 2 from both sides /2 /2 x = -8 would be correct ^_^